40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 15 01:41:43 2014, Jenne, Update, LSC, Data for DARM on sqrtInv investigation 
    Reply  Wed Jul 16 01:04:09 2014, Jenne, Update, LSC, Data for DARM on sqrtInv investigation 
Message ID: 10200     Entry time: Tue Jul 15 01:41:43 2014     Reply to this: 10208
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Data for DARM on sqrtInv investigation 

I took some data tonight for a quick look at what combinations of DC signals might be good to use for DARM, as an alternative to ALS before we're ready for RF.

I had the arms locked with ALS, PRMI with REFL33, and tried to move the CARM offset between plus and minus 1.  The PRMI wasn't holding lock closer than about -0.3 or +0.6, so that is also a problem.  Also, I realized just now that I have left the beam dumps in front of the transmission QPDs, so I had prevented any switching of the trans PD source.  This means that all of my data for C1:LSC-TR[x,y]_OUT_DQ is taken with the Thorlabs PDs, which is fine, although they saturate around arm powers of 4 ever since my analog gain increase on the whitening board.  Anyhow, the IFO didn't hold lock for much beyond then anyway, so I didn't miss out on much.  I need to remember to remove the dumps though!!

Self:  Good stuff should be between 12:50am - 1:09am.  One set of data was ./getdata -s 1089445700 -d 30 -c C1:LSC-TRX_OUT_DQ C1:LSC-TRY_OUT_DQ C1:LSC-CARM_IN1_DQ C1:LSC-PRCL_IN1_DQ

ELOG V3.1.3-